### v1

# BIBO stability of port-Hamiltonian systems

Abstract.
We study the question of when a distributed port-Hamiltonian system is bounded-input bounded-output (BIBO) stable. Exploiting the particular structure of the transfer function of these systems, we derive several sufficient conditions for BIBO stability.
*Corresponding author
Wierzba is supported by the Theme Team project “Predictive Avatar Control and Feedback” sponsored by the Faculty of Eletrical Engineering, Computer Science and Mathematics.
1. Introduction
Since their introduction port-Hamiltonian systems (pHS) have constituted an important system class used in modelling and control of real world applications. Their close connection to the ideas of energy-flow and energy-conservation makes them a useful tool in the study of physical systems [27, 28, 6].
A particular type of pHS are so-called distributed port-Hamiltonian systems [29, 13, 12, 30]. These contain systems formally given by a PDE of the form
$\frac{\partial x}{\partial t}=P_{1}\frac{\partial}{\partial\xi}\left(\mathcal{H}x\right)+P_{0}\left(\mathcal{H}x\right)$
on some interval $[a,b]$ together with boundary control and observation at its endpoints. The justification and motivation for the study of theses systems lies in their ability to provide a modelling framework for a variety of different real-world processes, including first of all transport and flow phenomena such as in fluid mechanics [17] or chemistry [34, 33] applications and also vibrations within flexible structures as they appear e.g. in robotics [15, 14]. In particular in this latter case, the approach using infinite-dimensional systems allows to produce refinements of finite-dimensional models where the mentioned phenomena (vibrations, flows, etc. ) are usually ignored or only treated in a coarse way. We refer the interested reader to the literature review [18] for a more extensive look on the history, study and application of distributed pHS.
Only recently work has been undertaken to systematically classify different notions of stability for distributed port-Hamiltonian systems. These studies considered in particular exponential [24] and asymptotic stability [32]. Previously only some sufficient conditions for these properties where known [12, 30], which were however mainly restricted to the case of systems featuring a contraction semigroup as encountered in the vast majority of concrete examples.
Besides these different internal stability concepts, the input-output behaviour of pHS is an important aspect. It is a nice feature that a distributed pHS – under practically weak technical assumptions – is $L^{2}$-well-posed [12]. Thus in particular there exist constants $c_{t}$ for any $t>0$ such that for any input $u\in L^{2}\left([0,t],U\right)$ the corresponding output $y$ satisfies $\|y\|_{L^{2}\left([0,t],Y\right)}\leq c_{t}\|u\|_{L^{2}\left([0,t],U\right)}$. The importance of this result is again grounded in the port-Hamiltonian structure being used to model energy-flows and the input-output stability thus being related to the concepts of energy-conserving and energy-dissipating systems.
In this contribution we study the notion of bounded-input bounded-output stability (BIBO stability), which is classical for finite-dimensional systems, for distributed port-Hamiltonian systems. An input-output system is called BIBO stable if any bounded input is mapped to a bounded output and there exists a uniform relative norm bound [26, 4, 31, 5, 1, 21, 9].
The question of whether a system is BIBO stable or not, while relatively straightforward in the finite-dimensional setting, becomes significantly more involved in the case of an infinite-dimensional state space such as for the pHS considered here. For example, it is then no longer the case that BIBO stability is always implied by exponential stability (see [21, Thm. 5.1] but note that crucially this example is not of port-Hamiltonian form).
In [21] it was shown that, for a linear system described in the framework of system nodes (see [23]) with finite-dimensional input and output spaces — as will be the case in the system class considered in this contribution — BIBO stability is equivalent to the inverse Laplace transform of the transfer function $\mathbf{G}(s)$ being a measure of bounded total variation. This includes in particular the case that the impulse response – upon being well-defined – of the system is an $L^{1}$ function.
While this condition offers a complete characterisation of when a system is BIBO stable, it is often of limited applicability for concrete systems. Typically no closed form expression for the inverse Laplace transform of the respective transfer function exists or is prohibitively hard to find. Thus BIBO stability is in general much more difficult to establish than e.g. $L^{2}$-well-posedness. This circumstance justifies the search for other sufficient or necessary conditions and the concrete study of BIBO stability for particular system classes. A few such results have been provided in for example [21] already, however almost exclusively for parabolic systems, which exclude pHSs as considered here.
We note that BIBO stability — besides being an intriguing topic in itself and classical control theory — plays an essential role in the applicability of funnel control to relative degree systems [10, 2]. Such systems with interior dynamics modelled by port-Hamiltonian systems appear for example in the study of mechanical systems containing flexible and deformable parts [15], which significantly influence the behaviour.
This paper begins with a recap of the definition of a class of distributed port-Hamiltonian systems in Section 2. Then, in Section 3, we derive an abstract decomposition of the transfer function and exploit its particular structure in order to apply the known characterisation of BIBO stability in a general setting. This yields a collection of sufficient criteria expressed in the form of matrix conditions for a subclass of pHS. Section 4 then presents a different approach to BIBO without any reference to transfer functions, instead employing the $L^{1}$-well-posedness of pHS on an $L^{1}$ state space and the connection between $L^{1}$ and $L^{\infty}$ stability notions discussed in [20].
1.1. Notation
For any $\alpha\in\mathbb{R}$, let $\mathbb{C}_{\alpha}:=\left\{z\in\mathbb{C}\;\middle|\;\mathrm{Re}\left(z\right)>\alpha\right\}$. Let $\mathbb{K}$ denote either $\mathbb{R}$ or $\mathbb{C}$. The Banach spaces considered are over the field $\mathbb{K}$.
Let $A$ be the generator of a strongly continuous semigroup $\mathbb{T}=\left(\mathbb{T}(t)\right)_{t\geq 0}$ on a Banach space $X$. $X_{1}$ is defined as the space $\mathrm{D}(A)$ with the norm $\|x\|_{X_{1}}:=\|(\beta I-A)x\|_{X}$ with $\beta\in\rho(A)$ and $X_{-1}$ as the completion of $X$ with respect to the norm $\|x\|_{X_{-1}}:=\|(\beta I-A)^{-1}x\|_{X}$ again with $\beta\in\rho(A)$. There exists a unique extension $\mathbb{T}_{-1}=\left(\mathbb{T}_{-1}(t)\right)_{t\geq 0}$ of the semigroup $\mathbb{T}$ to the space $X_{-1}$ with generator $A_{-1}:X\rightarrow X_{-1}$ which is an extension of the operator $A$. For more details see [23, 25].
For a matrix $M\in\mathbb{K}^{n\times n}$ and $p,q\in[1,\infty]$ let $\|M\|_{\ell^{p}\rightarrow\ell^{q}}$ be the matrix norm defined with respect to the $\ell^{p}$-norms on $\mathbb{K}^{n}$ given by $\|x\|_{\ell^{p}}=\left(\sum_{k=1}^{n}(x_{k})^{p}\right)^{\frac{1}{p}}$ for $p\in[1,\infty)$ and $\|x\|_{\ell^{\infty}}=\sup_{1\leq k\leq n}|x_{k}|$ for $p=\infty$ as
$\|M\|_{\ell^{p}\rightarrow\ell^{q}}=\sup_{x\in\mathbb{K}}\frac{\|Mx\|_{\ell^{q}}}{\|x\|_{\ell^{p}}}.$
For a Borel measure $h$ on $[0,\infty)$, let $\mathcal{L}\left\{h\right\}$ denote its Laplace transform if it exists (on some right half-plane). For a function $g$ of one complex variable, if there exists a measure $\mu$ such that $\mathcal{L}\left\{\mu\right\}$ and $g$ agree on some right half plane we will call $\mu$ the inverse Laplace transform and write $\mu=\mathcal{L}^{-1}\left\{g\right\}$.
Let $\mathcal{M}([0,\infty),\mathbb{K}^{n\times m})$ denote the set of Borel measures of bounded total variation on $[0,\infty)$ with values in $\mathbb{K}^{n\times m}$. Furthermore, for $h\in\mathcal{M}([0,\infty),\mathbb{K}^{n\times m})$ let $\|h\|_{\mathcal{M}}$ denote the total variation of $h$ [8, Sec. 3.2].
2. Distributed port-Hamiltonian systems
In this work we consider linear distributed port-Hamiltonian systems [12] of the form
(2.1) $\frac{\partial x}{\partial t}(\xi,t)$ $=P_{1}\frac{\partial}{\partial\xi}\left(\mathcal{H}(\xi)x(\xi,t)\right)+P_{0}(\xi)\left(\mathcal{H}(\xi)x(\xi,t)\right),$ $(\xi,t)$ $\in[a,b]\times[0,\infty),$
$x(\xi,0)$ $=x_{0}(\xi),$ $\xi$ $\in[a,b],$
$u(t)$ $=\widetilde{W}_{B}\begin{bmatrix}\left(\mathcal{H}x\right)(b,t)\\ \left(\mathcal{H}x\right)(a,t)\end{bmatrix},$ $t$ $\in[0,\infty),$
$y(t)$ $=\widetilde{W}_{C}\begin{bmatrix}\left(\mathcal{H}x\right)(b,t)\\ \left(\mathcal{H}x\right)(a,t)\end{bmatrix},$ $t$ $\in[0,\infty),$
with the following assumptions
  • $P_{1}\in\mathbb{K}^{n\times n}$ is invertible and self-adjoint;
  • $P_{0}\in L^{\infty}\left([a,b],\mathbb{K}^{n\times n}\right)$;
  • $\mathcal{H}\in L^{\infty}\left([a,b],\mathbb{K}^{n\times n}\right)$, $\mathcal{H}(\xi)$ is self-adjoint for almost all $\xi\in[a,b]$ and $\exists M,m>0$ s.t. $mI\leq\mathcal{H}(\xi)\leq MI$ for almost all $\xi\in[a,b]$;
  • $\widetilde{W}_{B}\in\mathbb{K}^{n\times 2n}$ and $\left[\begin{smallmatrix}\widetilde{W}_{B}\\ \widetilde{W}_{C}\end{smallmatrix}\right]\in\mathbb{K}^{2n\times 2n}$ have full rank.
We will in the following use $\textrm{pH}_{n}\left(P_{1},P_{0},\mathcal{H},\widetilde{W}_{B},\widetilde{W}_{C}\right)$ to denote this pHS.
In light of most of the pHS literature where contractive systems are considered, we explicitly state that we are making only the following weaker assumption.
Assumption 2.1.
The operator $A:=P_{1}\frac{\partial}{\partial\xi}\mathcal{H}+P_{0}\mathcal{H}$ with domain
$\mathrm{D}(A)=\left\{x_{0}\in X\,\middle|\,\mathcal{H}x_{0}\in H^{1}\left([a,b],\mathbb{K}^{n}\right),\widetilde{W}_{B}\begin{bmatrix}\left(\mathcal{H}x_{0}\right)(b)\\ \left(\mathcal{H}x_{0}\right)(a)\end{bmatrix}=0\right\}$
is the infinitesimal generator of a $C_{0}$-semigroup on the state space $X=L^{2}\left([a,b],\mathbb{K}^{n}\right)$ with inner product $\left\langle f,g\right\rangle_{X}:=\frac{1}{2}\int_{a}^{b}g(\xi)^{*}\mathcal{H}(\xi)f(\xi)\,\mathrm{d}\xi$.
Remark 2.2.
  1. (1)
    Note in particular, that we do not assume at this point that $A$ generates a contraction semigroup.
  2. (2)
    As the term $P_{0}\mathcal{H}$ represents merely a bounded perturbation within the operator $A$ that does not change its semigroup generation properties, Assumption 2.1 is in fact independent of the choice of $P_{0}$.
  3. (3)
    Contrary to the setting usually considered in the literature, we allow $P_{0}$ to depend on $\xi$. Note that the majority of results found e.g. in [12] for pHS with constant $P_{0}$ hold analogously in the case considered here.
  4. (4)
    The assumption of $A$ generating a $C_{0}$-semigroup is in particular satisfied if
    (2.2) $\widetilde{W}_{B}R_{0}^{-1}\Sigma\left(\widetilde{W}_{B}R_{0}^{-1}\right)^{*}\geq 0$
    where $R_{0}=\frac{1}{\sqrt{2}}\left[\begin{smallmatrix}P_{1}&-P_{1}\\ \mathbbm{1}_{n}&\mathbbm{1}_{n}\end{smallmatrix}\right]$ and $\Sigma=\left[\begin{smallmatrix}0&\mathbbm{1}_{n}\\ \mathbbm{1}_{n}&0\end{smallmatrix}\right]$. If in addition $P_{0}^{\ast}=-P_{0}$, then (2.2) is even equivalent to the semigroup being contractive [12, Thm. 7.2.4, Thm. 10.3.1].
For a pHS satisfying Assumption 2.1,direct extension of [12, Thm. 11.3.2 & Thm. 12.1.3] to the case of spatially dependent $P_{0}$, shows that its transfer function on $\rho(A)$ can be found as the unique solution of
(2.3) $\begin{split}&sx_{0}(\xi)=P_{1}\frac{\mathrm{d}}{\mathrm{d}\xi}\left(\mathcal{H}(\xi)x_{0}(\xi)\right)+P_{0}(\xi)\left(\mathcal{H}(\xi)x_{0}(\xi)\right),\\ &u_{0}=\widetilde{W}_{B}\begin{bmatrix}\left(\mathcal{H}x\right)(b)\\ \left(\mathcal{H}x\right)(a)\end{bmatrix},\qquad\mathbf{G}(s)u_{0}=\widetilde{W}_{C}\begin{bmatrix}\left(\mathcal{H}x\right)(b)\\ \left(\mathcal{H}x\right)(a)\end{bmatrix}.\end{split}$
2.1. BIBO stability of distributed pHS
In this paper we want to study the question of BIBO stability of such distributed port-Hamiltonian systems.
In the following we will understand the term classical solution in the common sense of evolution equations, that is as a collection $(u,x,y)$ with $u\in C([0,\infty),U)$, $x\in C^{1}([0,\infty),X)$ and $y\in C([0,\infty),Y)$ solving Equation (2.1).
Definition 2.3.
A pHS is called BIBO stable if there exists $c>0$ such that for any of its classical solution $(u,x,y)$ with $u\in C^{\infty}_{c}([0,\infty),U)$ and $x(0)=0$ and for all $t>0$ we have that $\|y\|_{L^{\infty}\left([0,t],Y\right)}\leq c\|u\|_{L^{\infty}\left([0,t],U\right)}$.
Now, as any pHS satisfying Assumption 2.1 can be equivalently expressed within the framework of system nodes (see [23] for a general introduction and [12, Ch. 11][19] for the rewriting of pHS respectively general boundary control systems in this form), they fall within the class of systems studied in [21]. Hence, by Theorem 3.4 therein, BIBO stability is equivalent to the transfer function $\mathbf{G}$ being the Laplace transform of a measure of bounded total variation.
Remark 2.4.
  1. (1)
    As any pHS can be rewritten as a system node and following the analogous construction in this formalism [23], we can define a generalised solution $(u,x,y)$ of a pHS for any input function $u\in L^{1}_{\mathrm{loc}}\left([0,\infty),U\right)$ and initial value $x_{0}\in X$.
    Here the state is given by $x(t)=\mathbb{T}(t)x_{0}+\int_{0}^{t}\mathbb{T}_{-1}\left(t-s\right)Bu(s)\,\mathrm{d}s\in X_{-1}$, where $\mathbb{T}$ is the semigroup generated by the operator $A$ from Assumption 2.1, and the output in the sense of distributions as $y(t)=\frac{\mathrm{d}^{2}}{\mathrm{d}t^{2}}\widetilde{W}_{C}\left[\begin{smallmatrix}\left(\mathcal{H}z\right)(b,t)\\ \left(\mathcal{H}z\right)(a,s)\end{smallmatrix}\right]$ where $z(t):=\int_{0}^{t}(t-s)x(t)\,\mathrm{d}s$.
  2. (2)
    By [21, Thm. 3.5], any generalised solution $(u,x,y)$ of a BIBO stable pHS with $u\in L^{\infty}_{\mathrm{loc}}\left([0,\infty),U\right)$ satisfies $y\in L^{\infty}_{\mathrm{loc}}\left([0,\infty),Y\right)$ (instead of only being a distribution) and for any $t>0$ we have that $\|y\|_{L^{\infty}\left([0,t],Y\right)}\leq c\|u\|_{L^{\infty}\left([0,t],U\right)}$, i.e. the inequality from Definition 2.3 extends to generalised solutions with bounded inputs.
The proposition below shows that the question of BIBO stability of pHS is not trivial.
Proposition 2.5.
  1. (1)
    There are BIBO stable pHS satisfying Assumption 2.1.
  2. (2)
    There are pHS satisfying Assumption 2.1, Equation (2.2) and $P_{0}^{*}=-P_{0}$ that are not BIBO stable.
Proof.
  1. (1)
    Take $\textrm{pH}_{1}\left(1,0,1,\left[\begin{smallmatrix}1&0\end{smallmatrix}\right],\left[\begin{smallmatrix}0&1\end{smallmatrix}\right]\right)$. Then the system is just the trivially BIBO stable transport equation with input $u$ and output $y$
    $\frac{\partial x}{\partial t}(\xi,t)$ $=\frac{\partial}{\partial\xi}\left(x(\xi,t)\right),$ $x(\xi,0)$ $=x_{0}(\xi),$
    $u(t)$ $=x(b,t),$ $y(t)$ $=x(a,t).$
  2. (2)
    Take $\textrm{pH}_{1}\left(1,0,1,\left[\begin{smallmatrix}1&-1\end{smallmatrix}\right],\left[\begin{smallmatrix}0&1\end{smallmatrix}\right]\right)$. Then $\widetilde{W}_{B}R_{0}^{-1}\Sigma\left(\widetilde{W}_{B}R_{0}^{-1}\right)^{*}=0$ and thus the system satisfies Assumption 2.1 by Remark 2.2 (and the semigroup is even contractive). On the other hand, the constant input $u(t)=1\in L^{\infty}\left([0,\infty),\mathbb{R}\right)$, yields the output $y(t)=\lfloor t\rfloor$ and hence the system is not BIBO stable.
As the pHS considered here feature finite-dimensional input and output spaces, Theorem 3.4 from [21] provides a necessary and sufficient condition for BIBO stability. That is the system is BIBO stable if and only if the transfer function $\mathbf{G}$ is the Laplace transform of a measure of bounded total variation on $[0,\infty)$.
While in theory this allows to check any pHS of this form for BIBO stability, in practice this condition is often not usable. Consider as an example the system from the proof of Proposition 2.5. Using Theorem 12.2.1 from [12] we find that its transfer function is $\mathbf{G}(s)=\frac{1}{e^{s}-1}$. Already for this simple function no nice closed form expression exists for its inverse Laplace transform, preventing us from a simple application of the criterion. As it turns out, this is the case generically, rendering this approach to proofing respectively disproofing BIBO stability to be of little use in practice if looking at a particular pHS. However, we will see that the transfer function and the properties of its inverse Laplace transform can — somewhat surprisingly — still be used to derive sufficient conditions for subclasses of these systems.
Finally, note that – similar as for exponential stability [24] – no condition only in terms of the matrices $P_{1}$ and $\widetilde{W}_{B}$ can fully classify BIBO stability for pHS.
Proposition 2.6.
There exists a BIBO stable $\textrm{pH}_{n}\left(P_{1},P_{0},\mathcal{H},\widetilde{W}_{B},\widetilde{W}_{C}\right)$ and a ~$\widetilde{\mathcal{H}}\in L^{\infty}\left([a,b],\mathbb{K}^{n\times n}\right)$ such that $\textrm{pH}_{n}\left(P_{1},P_{0},\mathcal{H},\widetilde{W}_{B},\widetilde{W}_{C}\right)$ is not BIBO stable.
Proof.
[…]
Change now the Hamiltonian to arrive at the system $\textrm{pH}_{2}\left(\mathbbm{1}_{2},0,\left[\begin{smallmatrix}\frac{1}{2}&0\\ 0&1\end{smallmatrix}\right],\widetilde{W}_{B},\widetilde{W}_{C}\right)$ Let then $u=\left[\begin{smallmatrix}u_{1}\\ u_{2}\end{smallmatrix}\right]$ be any smooth input such that $u_{1}(t)=0$ and $u_{2}(t)=(-1)^{t+1}$ for $t\in\mathbb{N}$. Then one can show that $y_{2}(t)=x_{2}(0,t)=\frac{1}{9}(-1)^{t+1}\left(3t+4-4\left(-\frac{1}{2}\right)^{t+1}\right)$ for $t\in\mathbb{N}$. Thus $\lim_{t\rightarrow\infty}y_{2}(t)=\infty$ and the system cannot be BIBO stable. ∎
3. An approach to BIBO stability via decomposition of the transfer function
A first approach to deriving sufficient conditions for BIBO stability is based on applying the characterisation from [21] not in an individual case-by-case manner — which as we saw may be prohibitively hard — but instead to a general expression of the transfer function.
3.1. An expression for the transfer function
We begin by deriving a particular decomposition of the transfer function of a pHS as described above by methods inspired by the constructions carried out in [24] and [32] for the study of exponential respectively asymptotic stability.
For any pHS and $s\in\mathbb{C}$ we have the fundamental solution $\Psi_{\zeta}^{s}\in\mathbb{K}^{n\times n}$ of the associated ODE
(3.1) $v^{\prime}(\zeta)=-P_{1}^{-1}\left(P_{0}(\zeta)-s\mathcal{H}^{-1}(\zeta)\right)v(\zeta).$
on $[a,b]$ with $\Psi_{a}^{s}=\mathbbm{1}_{n}$ as a continuous function $\zeta\mapsto\Psi_{\zeta}^{s}$ due to [32, Cor. A.2].
Proposition 3.1.
The transfer function $\mathbf{G}(s)$ of $\textrm{pH}_{n}\left(P_{1},P_{0},\mathcal{H},\widetilde{W}_{B},\widetilde{W}_{C}\right)$ satisfying Assumption 2.1 is given for $s\in\rho(A)$ by
(3.2) $\begin{split}\mathbf{G}(s)=\widetilde{W}_{C}\begin{bmatrix}\Psi_{b}^{s}\\ \mathbbm{1}_{n}\end{bmatrix}\left(\widetilde{W}_{B}\begin{bmatrix}\Psi_{b}^{s}\\ \mathbbm{1}_{n}\end{bmatrix}\right)^{-1}.\end{split}$
Proof.
We can write the solution of the first equation from (LABEL:eq:transferFunctionODE) using $\Psi_{\zeta}^{s}$ as $x_{0}(\zeta)=\Psi_{\zeta}^{s}\left(\mathcal{H}x_{0}\right)(a)$, so that the boundary conditions become
$u_{0}=\widetilde{W}_{B}\begin{bmatrix}\Psi_{b}^{s}\\ \mathbbm{1}_{n}\end{bmatrix}\left(\mathcal{H}x\right)(a),\qquad\mathbf{G}(s)u_{0}=\widetilde{W}_{C}\begin{bmatrix}\Psi_{b}^{s}\\ \mathbbm{1}_{n}\end{bmatrix}\left(\mathcal{H}x\right)(a),$
from which the claim follows by solving for $\mathbf{G}(s)$. ∎
Remark 3.2.
Note that for a system with constant coefficients the fundamental solution is given by
(3.3) $\Psi_{\zeta}^{s}=e^{-P_{1}^{-1}\left(P_{0}-s\mathcal{H}^{-1}\right)(\zeta-a)}.$
For a non-constant Hamiltonian for which $P_{1}^{-1}\left(P_{0}(\xi)-s\mathcal{H}^{-1}(\xi)\right)$ commutes with $\int_{a}^{\zeta}P_{1}^{-1}\left(P_{0}(\tau)-s\mathcal{H}^{-1}(\tau)\right)\,\mathrm{d}\tau$ for all $\xi,\zeta\in[a,b]$ we have that
(3.4) $\Psi_{\zeta}^{s}=e^{-P_{1}^{-1}\int_{a}^{\zeta}\left(P_{0}-s\mathcal{H}^{-1}(\tau)\right)\,\mathrm{d}\tau}.$
3.2. Decomposing $\widetilde{W}_{B}$
As $P_{1}$ is self-adjoint and invertible we have the decomposition $\mathbb{K}^{n}=E_{+}\oplus E_{-}$ with
$E_{+}:=\textrm{lin}\left\{v\in\mathbb{K}^{n}\middle|\exists\lambda>0:P_{1}v=\lambda v\right\},\quad E_{-}:=\textrm{lin}\left\{v\in\mathbb{K}^{n}\middle|\exists\lambda<0:P_{1}v=\lambda v\right\}.$
Let $\iota_{\pm}:E_{\pm}\rightarrow\mathbb{K}^{n}$ denote the canonical embeddings and define
$P_{1}^{\pm}:=\iota_{\pm}^{*}\left(\pm P_{1}\right)\iota_{\pm}:E_{\pm}\rightarrow E_{\pm},\qquad Q_{\pm}:=\iota_{\pm}\left(P^{\pm}_{1}\right)^{\frac{1}{2}}\iota^{*}_{\pm}:\mathbb{K}^{n}\rightarrow\mathbb{K}^{n},$
where the second definition is possible as $P_{1}^{\pm}$ are both strictly positive self-adjoint. Note that, as $P_{\pm}:=\iota_{\pm}\iota_{\pm}^{*}:\mathbb{K}^{n}\rightarrow\mathbb{K}^{n}$ are the orthogonal projections on $E_{\pm}$ and since $E_{\pm}$ are invariant under $P_{1}$, we have that
$P_{1}=Q_{+}Q_{+}-Q_{-}Q_{-}\qquad\textrm{and}\qquad\mathbbm{1}_{n}=Q_{+}P_{1}^{-1}Q_{+}-Q_{-}P_{1}^{-1}Q_{-}.$
For the further study of the transfer function as expressed in Equation (3.2), it will be beneficial if the input matrix $\widetilde{W}_{B}$ can be decomposed in a particular way with respect to the matrices $Q_{+}$ and $Q_{-}$. We believe the following result providing this decomposition to be of interest in its own right.
Proposition 3.3.
  1. (1)
    There exist unique matrices $J,L\in\mathbb{K}^{n\times n}$ such that
    (3.5) $\widetilde{W}_{B}=\begin{pmatrix}JQ_{+}-LQ_{-}&JQ_{-}-LQ_{+}\end{pmatrix}.$
    In that case $J=\widetilde{W}_{B}\begin{bmatrix}Q_{+}\\ Q_{-}\end{bmatrix}P_{1}^{-1}$ and $L=\widetilde{W}_{B}\begin{bmatrix}Q_{-}\\ Q_{+}\end{bmatrix}P_{1}^{-1}$.
  2. (2)
    The following statements are equivalent:
    1. (a)
      There exist matrices $K,M\in\mathbb{K}^{n\times n}$ with $K$ invertible such that
      (3.6) $\widetilde{W}_{B}=K\begin{pmatrix}Q_{+}-MQ_{-}&Q_{-}-MQ_{+}\end{pmatrix}.$
    2. (b)
      The matrix $\widetilde{W}_{B}\begin{bmatrix}P_{+}\\ P_{-}\end{bmatrix}$ is invertible.
    3. (c)
      The operator $A_{\mathbbm{1}}:=P_{1}\frac{\partial}{\partial\xi}+P_{0}$ with domain
      $\mathrm{D}(A_{\mathbbm{1}})=\left\{x_{0}\in X\,\middle|\,x_{0}\in H^{1}\left([a,b],\mathbb{K}^{n}\right),\widetilde{W}_{B}\begin{bmatrix}x_{0}(b)\\ x_{0}(a)\end{bmatrix}=0\right\}$
      is the infinitesimal generator of a $C_{0}$-semigroup on $L^{2}\left([a,b],\mathbb{K}^{n}\right)$.
    4. (d)
      The matrix $J$ from Part 1 is invertible.
    Furthermore, in this case $K=\widetilde{W}_{B}\begin{bmatrix}Q_{+}\\ Q_{-}\end{bmatrix}P_{1}^{-1}$ and $M=K^{-1}\widetilde{W}_{B}\begin{bmatrix}Q_{-}\\ Q_{+}\end{bmatrix}P_{1}^{-1}$.
Proof.
  1. (1)
    Let $\widetilde{W}_{B}=\begin{pmatrix}W_{1}&W_{2}\end{pmatrix}$ and define
    $J$ $:=\widetilde{W}_{B}\begin{bmatrix}Q_{+}\\ Q_{-}\end{bmatrix}P_{1}^{-1}=\left(W_{1}Q_{+}+W_{2}Q_{-}\right)P_{1}^{-1}$
    $L$ $:=\widetilde{W}_{B}\begin{bmatrix}Q_{-}\\ Q_{+}\end{bmatrix}P_{1}^{-1}=\left(W_{1}Q_{-}+W_{2}Q_{+}\right)P_{1}^{-1}.$
    Then $JQ_{+}-LQ_{-}=W_{1}\left(Q_{+}P_{1}^{-1}Q_{+}-Q_{-}P_{1}^{-1}Q_{-}\right)=W_{1}$ and similarly $JQ_{-}-LQ_{+}=W_{2}$, so that $\widetilde{W}_{B}=\begin{pmatrix}JQ_{+}-LQ_{-}&JQ_{-}-LQ_{+}\end{pmatrix}$.
    Conversely, assume that $\widetilde{W}_{B}$ is given in the form (3.5). Then
    $\widetilde{W}_{B}\begin{bmatrix}Q_{+}\\ Q_{-}\end{bmatrix}P_{1}^{-1}=\left(JQ_{+}-LQ_{-}\right)Q_{+}P_{1}^{-1}+\left(JQ_{-}-LQ_{+}\right)Q_{-}P_{1}^{-1}=J$
    and analogously for $L$, which shows the uniqueness.
  2. (2)
    By Part 1 of the Lemma and the definitions of $P_{\pm}$ and $Q_{\pm}$, clearly the decomposition (3.6) exists if and only if
    $J=\widetilde{W}_{B}\begin{bmatrix}Q_{+}\\ Q_{-}\end{bmatrix}P_{1}^{-1}=\widetilde{W}_{B}\begin{bmatrix}P_{+}\\ P_{-}\end{bmatrix}\left(Q_{+}+Q_{-}\right)P_{1}^{-1}$
    is invertible, giving the equivalence of (a) and (d). Further, as both $Q_{+}+Q_{-}$ and $P_{1}^{-1}$ are invertible, this is the case if and only if (b) holds.
    The equivalence of (b) and (c) is a consequence of Theorem 1.5 in [11].
    Finally, the expressions for $K$ and $M$ follow then directly from Part (1).
Remark 3.4.
  1. (1)
    Note that statement (c) of Proposition 3.3 could equivalently be formulated as $\textrm{pH}_{n}\left(P_{1},P_{0},\mathbbm{1}_{n},\widetilde{W}_{B},\widetilde{W}_{C}\right)$ satisfying Assumption 2.1.
  2. (2)
    A direct corollary of Proposition 3.3 is that a decomposition (3.6) exists in particular if $\widetilde{W}_{B}$ satisfies Equation (2.2), as in this case the semigroup generation property is independent of the Hamiltonian. This particular result was first derived by explicit calculation in [24, Lem. 2.3].
  3. (3)
    If a decomposition (3.6) holds true, then by direct calculation we find that
    $\widetilde{W}_{B}R_{0}^{-1}\Sigma\left(\widetilde{W}_{B}R_{0}^{-1}\right)^{*}=K\left(\mathbbm{1}_{n}-MM^{\ast}\right)K^{*},$
    so that in particular Equation (2.2) is equivalent to $\mathbbm{1}_{n}\geq MM^{\ast}$ respectively $\|M\|_{\ell^{2}\rightarrow\ell^{2}}\leq 1$. This is the second half of [24, Lem. 2.3].
  4. (4)
    An example of a pHS where no decomposition (3.6) exists is given by $P_{1}=\left[\begin{smallmatrix}0&1\\ 1&0\end{smallmatrix}\right]$, $P_{0}=0$, $\mathcal{H}(\xi)=\left[\begin{smallmatrix}1+\xi&0\\ 0&1\end{smallmatrix}\right]$ for $\xi\in[0,1]$ and $\widetilde{W}_{B}=\left[\begin{smallmatrix}1&0&-1&0\\ 0&1&0&1\end{smallmatrix}\right]$. Then by [11, Ex. 3.2] the operator $A_{\mathbbm{1}}$ does not generate a $C_{0}$-semigroup and thus by Proposition 3.3 no such decomposition exists.
For the remainder of this subsection we now take the assumption that the equivalent statements from Proposition 3.3 hold true.
Assumption 3.5.
There exist matrices $K,M\in\mathbb{K}^{n\times n}$ with $K$ invertible such that
(3.7) $\widetilde{W}_{B}=K\begin{pmatrix}Q_{+}-MQ_{-}&Q_{-}-MQ_{+}\end{pmatrix}.$
We now consider the middle of the three parts from the expression for $\mathbf{G}(s)$ in Equation (3.2) closer. Note that with the decomposition (3.7) we have
$\widetilde{W}_{B}\begin{bmatrix}\Psi_{b}^{s}\\ \mathbbm{1}_{n}\end{bmatrix}$ $=K\begin{pmatrix}Q_{+}-MQ_{-}&Q_{-}-MQ_{+}\end{pmatrix}\begin{bmatrix}\Psi_{b}^{s}\\ \mathbbm{1}_{n}\end{bmatrix}$
$=K\left(Q_{-}+Q_{+}\Psi_{b}^{s}\right)-KM\left(Q_{+}+Q_{-}\Psi_{b}^{s}\right).$
Lemma 3.6.
The matrix $Q_{-}+Q_{+}\Psi_{b}^{s}$ is invertible for any $s\in\rho(\widetilde{A})$ where $\widetilde{A}$ is the operator associated to $\textrm{pH}_{n}\left(P_{1},P_{0},\mathcal{H},\left[\begin{smallmatrix}Q_{+}&Q_{-}\end{smallmatrix}\right],\widetilde{W}_{C}\right)$. In particular there exists $\alpha\in\mathbb{R}$ such that $\mathbb{C}_{\alpha}\subset\rho(\widetilde{A})$ and hence $Q_{-}+Q_{+}\Psi_{b}^{s}$ is invertible for all $s\in\mathbb{C}_{\alpha}$.
Proof.
Note first that $\widetilde{A}$ indeed satisfies Assumption 2.1, as it generates a $C_{0}$-semigroup by [12, Thm. 7.2.4, Thm. 10.3.1] as
$\begin{bmatrix}Q_{+}&Q_{-}\end{bmatrix}R_{0}^{-1}\Sigma\left(\begin{bmatrix}Q_{+}&Q_{-}\end{bmatrix}R_{0}^{-1}\right)^{*}=Q_{+}P_{1}^{-1}Q_{+}-Q_{-}P_{1}^{-1}Q_{-}=\mathbbm{1}_{n}>0.$
By Proposition 3.1 the transfer function of the system $\textrm{pH}_{n}\left(P_{1},P_{0},\mathcal{H},\left[\begin{smallmatrix}Q_{+}&Q_{-}\end{smallmatrix}\right],\widetilde{W}_{C}\right)$ can thus be represented on $\rho(\widetilde{A})$ as $\widetilde{\mathbf{G}}(s)=\widetilde{W}_{C}\left[\begin{smallmatrix}\widetilde{\Psi}_{b}^{s}\\ \mathbbm{1}_{n}\end{smallmatrix}\right]\left(\left[\begin{smallmatrix}Q_{+}&Q_{-}\end{smallmatrix}\right]\left[\begin{smallmatrix}\widetilde{\Psi}_{b}^{s}\\ \mathbbm{1}_{n}\end{smallmatrix}\right]\right)^{-1}$ where $\widetilde{\Psi}_{\zeta}^{s}$ is the fundamental solution associated to that system. Hence in particular $\left[\begin{smallmatrix}Q_{+}&Q_{-}\end{smallmatrix}\right]\left[\begin{smallmatrix}\widetilde{\Psi}_{b}^{s}\\ \mathbbm{1}_{n}\end{smallmatrix}\right]=Q_{-}+Q_{+}\widetilde{\Psi}_{b}^{s}$ is invertible. But from the definition of the fundamental solution it is clear that $\widetilde{\Psi}_{\zeta}^{s}=\Psi_{\zeta}^{s}$ thus proving the claim. ∎
Remark 3.7.
Under the additional assumptions that $\mathcal{H}\in C^{1}\left([a,b],\mathbb{K}^{n\times n}\right)$ and $P_{0}=-P_{0}^{*}$, the operator $\widetilde{A}$ even generates an exponentially stable contraction semigroup [12, Lem. 9.1.4]. Then there even exists a negative $\alpha<0$ satisfying the conditions in Lemma 3.6. Thus $Q_{-}+Q_{+}\Psi_{b}^{s}$ and $(Q_{-}+Q_{+}\Psi_{b}^{s})^{-1}$ are then in particular defined on the imaginary axis and the whole positive right half plane.
Definition 3.8.
For $s\in\mathbb{C}_{\alpha}$ with $\alpha$ from Lemma 3.6, define the matrices
$\mathbf{V}(s):=Q_{-}+Q_{+}\Psi_{b}^{s}\qquad\textrm{and}\qquad\mathbf{U}(s):=\left(Q_{+}+Q_{-}\Psi_{b}^{s}\right)\mathbf{V}(s)^{-1}.$
Using these matrices we can now simplify the expression for the transfer function.
Lemma 3.9.
The transfer function of a pHS satisfying Assumptions 2.1 and 3.5 is given on $s\in\rho(A)\cap\mathbb{C}_{\alpha}$ with $\alpha$ from Lemma 3.6 by $\mathbf{G}(s)=\mathbf{Z}(s)\left(1-M\mathbf{U}(s)\right)^{-1}K^{-1},$ where $\mathbf{Z}(s)=\widetilde{W}_{C}\begin{bmatrix}\Psi_{b}^{s}\\ \mathbbm{1}_{n}\end{bmatrix}\mathbf{V}(s)^{-1}$. This holds in particular on $\mathbb{C}_{\widetilde{\alpha}}\subset\rho(A)\cap\mathbb{C}_{\alpha}$ for some $\widetilde{\alpha}\geq\alpha$.
Proof.
This follows from direct calculation, as
$\mathbf{G}(s)$ $=\widetilde{W}_{C}\begin{bmatrix}\Psi_{b}^{s}\\ \mathbbm{1}_{n}\end{bmatrix}\left(\widetilde{W}_{B}\begin{bmatrix}\Psi_{b}^{s}\\ \mathbbm{1}_{n}\end{bmatrix}\right)^{-1}=\widetilde{W}_{C}\begin{bmatrix}\Psi_{b}^{s}\\ \mathbbm{1}_{n}\end{bmatrix}\left(K\mathbf{V}(s)-KM\mathbf{U}(s)\mathbf{V}(s)\right)^{-1}$
$=\widetilde{W}_{C}\begin{bmatrix}\Psi_{b}^{s}\\ \mathbbm{1}_{n}\end{bmatrix}\mathbf{V}(s)^{-1}\left(\mathbbm{1}_{n}-M\mathbf{U}(s)\right)^{-1}K^{-1}=\mathbf{Z}(s)\left(\mathbbm{1}_{n}-M\mathbf{U}(s)\right)^{-1}K^{-1}.\qed$
As mentioned before, one problem in deciding the question of BIBO stability using the characterisation from [21] is finding the inverse Laplace transform of $\mathbf{G}$. Even in the case of a constant Hamiltonian where the fundamental solution $\Psi_{\zeta}^{s}$ takes the simple form (3.3) of a matrix exponential, we see that due to the inverse $\left(\widetilde{W}_{B}\left[\begin{smallmatrix}\Psi_{b}^{s}\\ \mathbbm{1}_{n}\end{smallmatrix}\right]\right)^{-1}$ it will in general be difficult to find that inverse transform. With the representation from Lemma 3.9 however, this problem can be circumnavigated in some cases and lead to a sufficient condition for BIBO stability.
Lemma 3.10.
  1. (1)
    Let $s\in\mathbb{C}_{\widetilde{\alpha}}$ with $\widetilde{\alpha}$ from Lemma 3.9. Then if the series $\sum_{k=0}^{\infty}\left(M\mathbf{U}(s)\right)^{k}$ converges we have that
    (3.8) $\mathbf{G}(s)=\mathbf{Z}(s)\left(\sum_{k=0}^{\infty}\left(M\mathbf{U}(s)\right)^{k}\right)K^{-1}.$
  2. (2)
    If $\mathcal{L}^{-1}\left\{\mathbf{Z}\right\}$ and $\mathcal{L}^{-1}\left\{\left(M\mathbf{U}\right)^{k}\right\}$ exist as measures of bounded total variation for all $k>0$ and in addition
    $\sum_{k=0}^{\infty}\left\|\mathcal{L}^{-1}\left\{\left(M\mathbf{U}\right)^{k}\right\}\right\|_{\mathcal{M}}<\infty,$
    then $\textrm{pH}_{n}\left(P_{1},P_{0},\mathcal{H},\widetilde{W}_{B},\widetilde{W}_{C}\right)$ is BIBO stable and we have
    (3.9) $\mathcal{L}^{-1}\left\{\mathbf{G}\right\}=\mathcal{L}^{-1}\left\{\mathbf{Z}\right\}\ast\left(\sum_{k=0}^{\infty}\mathcal{L}^{-1}\left\{\left(M\mathbf{U}\right)^{k}\right\}\right)K^{-1}.$
Proof.
  1. (1)
    Clear.
  2. (2)
    By the assumptions the right hand side of Equation (3.9) is well-defined as a measure of bounded total variation $h\in\mathcal{M}\left([0,\infty),\mathbb{K}^{n\times n}\right)$.
    Term-by-term Laplace transforming of the right hand side of Equation (3.9) yields the right hand side of Equation (3.8) and the continuity of the Laplace transform implies that this sum indeed converges to $\mathcal{L}\left\{h\right\}$ or equivalently that Equation (3.9) holds. BIBO stability then follows from [21, Thm. 3.4]. ∎
3.3. Equivalent diagonal system
For the rest of this section we will now take the following additional assumption.
Assumption 3.11.
There exist $S,\mathcal{H}^{D}\in C^{1}\left([a,b],\mathbb{K}^{n\times n}\right)$ with $S$ pointwise invertible, $\mathcal{H}^{D}$ diagonal and positive such that $P_{1}\mathcal{H}=S^{-1}P^{D}_{1}\mathcal{H}^{D}S$, where
$P_{1}^{D}=\begin{pmatrix}\mathbbm{1}_{m}&0\\ 0&-\mathbbm{1}_{n-m}\end{pmatrix}$
for some $0\leq m\leq n$.
Remark 3.12.
  1. (1)
    It is a well-know result that any pHS satisfying Assumption 3.11 is regular and $L^{2}$-well-posed [12, Thm. 13.2.2].
  2. (2)
    Using the same splitting as in $P_{1}^{D}$ we can write $\mathcal{H}^{D}=\begin{pmatrix}\mathcal{H}^{D}_{+}&0\\ 0&\mathcal{H}^{D}_{-}\end{pmatrix}.$ Then +$\mathcal{H}^{D}_{+}$ contains the positive and $-\mathcal{H}^{D}_{-}$ the negative eigenvalues of $P_{1}\mathcal{H}$ on the respective diagonals. Furthermore the matrix $S^{-1}(\xi)$ then has the form
    $S^{-1}(\xi)=\begin{pmatrix}e_{1}(\xi)&\cdots&e_{m}(\xi)&e_{m+1}(\xi)&\cdots&e_{n}(\xi)\end{pmatrix}$
    with $e_{1}(\xi),\ldots,e_{m}(\xi)$ eigenvectors of $P_{1}\mathcal{H}(\xi)$ corresponding to the eigenvalues in +$\mathcal{H}^{D}_{+}(\xi)$ and $e_{m+1}(\xi),\ldots,e_{n}(\xi)$ those corresponding to the ones in $-\mathcal{H}^{D}_{-}(\xi)$.
We now consider the state transformation $\widetilde{x}=Sx$. Then the PDE in Equation (2.1) describing the port-Hamiltonian system becomes
$\frac{\partial\widetilde{x}}{\partial t}(\xi,t)$ $=P_{1}^{D}\frac{\partial}{\partial\xi}\left(\mathcal{H}^{D}(\xi)\widetilde{x}(\xi,t)\right)+P_{0}^{D}(\xi)\left(\mathcal{H}^{D}(\xi)\widetilde{x}(\xi,t)\right),$
with $P_{0}^{D}(\xi):=S(\xi)\left[\frac{\mathrm{d}S^{-1}(\xi)}{\mathrm{d}\xi}+P_{0}P_{1}^{-1}S^{-1}(\xi)\right]P_{1}^{D}$.
Similarly, the boundary control and observation also change under the state transformation and become
$u(t)$ $=\widetilde{W}_{B}\begin{bmatrix}P_{1}^{-1}S^{-1}(b)P_{1}^{D}&0\\ 0&P_{1}^{-1}S^{-1}(a)P_{1}^{D}\end{bmatrix}\begin{bmatrix}\left(\mathcal{H}^{D}\widetilde{x}\right)(b,t)\\ \left(\mathcal{H}^{D}\widetilde{x}\right)(a,t)\end{bmatrix}=\widetilde{W}_{B}^{D}\begin{bmatrix}\left(\mathcal{H}^{D}\widetilde{x}\right)(b,t)\\ \left(\mathcal{H}^{D}\widetilde{x}\right)(a,t)\end{bmatrix},$
$y(t)$ $=\widetilde{W}_{C}\begin{bmatrix}P_{1}^{-1}S^{-1}(b)P_{1}^{D}&0\\ 0&P_{1}^{-1}S^{-1}(a)P_{1}^{D}\end{bmatrix}\begin{bmatrix}\left(\mathcal{H}^{D}\widetilde{x}\right)(b,t)\\ \left(\mathcal{H}^{D}\widetilde{x}\right)(a,t)\end{bmatrix}=\widetilde{W}_{C}^{D}\begin{bmatrix}\left(\mathcal{H}^{D}\widetilde{x}\right)(b,t)\\ \left(\mathcal{H}^{D}\widetilde{x}\right)(a,t)\end{bmatrix}.$
Hence we find that the diagonalised system is just $\textrm{pH}_{n}\left(P_{1}^{D},P_{0}^{D},\mathcal{H}_{D},\widetilde{W}_{B}^{D},\widetilde{W}_{C}^{D}\right)$. As the system resulting from the state transformation has the same input-output behaviour as the original system, they have in particular the same transfer function and thus one is BIBO stable if and only if the other one is.
However, applying the results from the previous section is much simpler in the diagonalised system due to the special structure of $P_{1}^{D}$. We can for example easily calculate the associated matrices $Q^{D}_{+}=\left[\begin{smallmatrix}\mathbbm{1}_{m}&0\\ 0&0\end{smallmatrix}\right]$ and $Q^{D}_{-}=\left[\begin{smallmatrix}0&0\\ 0&\mathbbm{1}_{n-m}\end{smallmatrix}\right]$ used in the expression for the transfer function derived earlier.
Furthermore we find that the transformed input matrix $\widetilde{W}_{B}^{D}$ of the diagonalised system always possesses a decomposition of the form (3.7).
Proposition 3.13.
The matrix $\widetilde{W}^{D}_{B}\begin{bmatrix}P^{D}_{+}\\ P^{D}_{-}\end{bmatrix}$ is invertible and hence there exist matrices $K,M\in\mathbb{K}^{n\times n}$ with $K$ invertible such that
$\widetilde{W}_{B}^{D}=K\begin{pmatrix}Q_{+}^{D}-MQ_{-}^{D}&Q_{-}^{D}-MQ_{+}^{D}\end{pmatrix}.$
Proof.
Using that $P_{1}\mathcal{H}=S^{-1}P_{1}^{D}\mathcal{H}^{D}S$ we find that
$\widetilde{W}_{B}^{D}$ $=\widetilde{W}_{B}\begin{bmatrix}P_{1}^{-1}S^{-1}(b)P_{1}^{D}&0\\ 0&P_{1}^{-1}S^{-1}(a)P_{1}^{D}\end{bmatrix}$
$=\widetilde{W}_{B}\begin{bmatrix}\mathcal{H}(b)S^{-1}(b)\left(\mathcal{H}^{D}\right)^{-1}(b)&0\\ 0&\mathcal{H}(a)S^{-1}(a)\left(\mathcal{H}^{D}\right)^{-1}(a)\end{bmatrix}.$
Letting $\widetilde{W}_{B}=\begin{pmatrix}W_{1}&W_{2}\end{pmatrix}$ we then have
$\widetilde{W}^{D}_{B}\begin{bmatrix}P^{D}_{+}\\ P^{D}_{-}\end{bmatrix}$ $=\widetilde{W}_{B}\begin{bmatrix}\mathcal{H}(b)S^{-1}(b)\left(\mathcal{H}^{D}\right)^{-1}(b)P^{D}_{+}\\ \mathcal{H}(a)S^{-1}(a)\left(\mathcal{H}^{D}\right)^{-1}(a)P^{D}_{-}\end{bmatrix}$
$=W_{1}\mathcal{H}(b)S^{-1}(b)\left(\mathcal{H}^{D}\right)^{-1}(b)P^{D}_{+}+W_{2}\mathcal{H}(a)S^{-1}(a)\left(\mathcal{H}^{D}\right)^{-1}(a)P^{D}_{-}.$
Considering now the form of $S^{-1}$, $\mathcal{H}^{D}$ and $P_{\pm}^{D}$, we see that
$S^{-1}(b)\left(\mathcal{H}^{D}\right)^{-1}(b)P^{D}_{+}$ $=\begin{pmatrix}\frac{e_{1}(b)}{\lambda_{1}(b)}&\cdots&\frac{e_{m}(b)}{\lambda_{m}(b)}&0&\cdots&0\end{pmatrix}$
$S^{-1}(a)\left(\mathcal{H}^{D}\right)^{-1}(a)P^{D}_{-}$ $=\begin{pmatrix}0&\cdots&0&\frac{e_{m+1}(a)}{\lambda_{m+1}(a)}&\cdots&\frac{e_{n}(a)}{\lambda_{n}(a)}&\end{pmatrix},$
where $\lambda_{i}(\xi)$ are the entries of ±$\mathcal{H}^{D}_{\pm}(\xi)$ corresponding to the eigenvectors $e_{i}(\xi)$ of $P_{1}\mathcal{H}(\xi)$.
We thus realise that the range of $\widetilde{W}^{D}_{B}\begin{bmatrix}P^{D}_{+}\\ P^{D}_{-}\end{bmatrix}$ is given by
$W_{1}\,\mathcal{H}(b)\,\textrm{span}\left\{e_{i\leq m}(b)\right\}\oplus W_{2}\,\mathcal{H}(a)\,\textrm{span}\left\{e_{i>m}(a)\right\}.$
But by Theorem 1.5 in [11] this space is all of $\mathbb{K}^{n}$ as the operator $A$ generates a $C_{0}$-semigroup by Assumption 2.1. Hence $\widetilde{W}^{D}_{B}\begin{bmatrix}P^{D}_{+}\\ P^{D}_{-}\end{bmatrix}$ is invertible and existence of the decomposition follows from Proposition 3.3. ∎
Remark 3.14.
  1. (1)
    Using the explicit expressions from Proposition 3.3, we find that the matrix $M$ from Proposition 3.13 is given by
    $M=\left(\widetilde{W}_{B}\begin{bmatrix}P_{1}^{-1}S^{-1}(b)P^{D}_{+}\\ P_{1}^{-1}S^{-1}(a)P^{D}_{-}\end{bmatrix}\right)^{-1}\widetilde{W}_{B}\begin{bmatrix}P_{1}^{-1}S^{-1}(b)P^{D}_{-}\\ P_{1}^{-1}S^{-1}(a)P^{D}_{+}\end{bmatrix}.$
  2. (2)
    Note that the proof of Proposition 3.13 shows that the existence of the decomposition of $\widetilde{W}^{D}_{B}$ is in fact equivalent to Assumption 2.1.
3.4. The case $P^{D}_{0}=0$
We now want to restrict ourselves to a particular type of pHS and thus take the following assumption for the rest of this section.
Assumption 3.15.
$P_{0}^{D}=0$
Remark 3.16.
  1. (1)
    Note that Assumption 3.15 is satisfied in particular if the pHS has vanishing $P_{0}=0$ and a Hamiltonian $\mathcal{H}$ that is constant along $[a,b]$.
  2. (2)
    There are cases in which $P_{0}^{D}=0$ even though the Hamiltonian is not constant and $P_{0}\neq 0$. In fact, $P_{0}^{D}$ vanishes if and only if $S^{-1}$ satisfies the differential equation $\frac{\mathrm{d}S^{-1}(\xi)}{\mathrm{d}\xi}=-P_{0}P_{1}^{-1}S^{-1}(\xi).$
    A concrete example on $[a,b]=[0,1]$ where this is the case despite non-constant $\mathcal{H}$ and $P_{0}\neq 0$ is given by $\textrm{pH}_{2}\left(\mathbbm{1}_{2},\left[\begin{smallmatrix}0&1\\ -1&0\end{smallmatrix}\right],\mathcal{H},\widetilde{W}_{B},\widetilde{W}_{C}\right)$ with
    $\mathcal{H}(\xi)=\frac{1}{2}\begin{bmatrix}2+3\xi-\xi\cos(2\xi)&-\xi\sin(2\xi)\\ -\xi\sin(2\xi)&2+3\xi+\xi\cos(2\xi)\end{bmatrix}.$
    Then we find $S^{-1}(\xi)=\left[\begin{smallmatrix}\cos(\xi)&-\sin(\xi)\\ \sin(\xi)&\cos(\xi)\end{smallmatrix}\right]$, $\mathcal{H}^{D}(\xi)=\left[\begin{smallmatrix}1+\xi&0\\ 0&1+2\xi\end{smallmatrix}\right]$ and $P_{0}^{D}=0$.
The reason for making this assumption is that in this case the diagonalised system takes the simple form
$\frac{\partial\widetilde{x}}{\partial t}(\xi,t)$ $=P_{1}^{D}\frac{\partial}{\partial\xi}\left(\mathcal{H}^{D}\widetilde{x}(\xi,t)\right),$
$u(t)$ $=\widetilde{W}_{B}^{D}\begin{bmatrix}\left(\mathcal{H}^{D}\widetilde{x}\right)(b,t)\\ \left(\mathcal{H}^{D}\widetilde{x}\right)(a,t)\end{bmatrix},$
$y(t)$ $=\widetilde{W}_{C}^{D}\begin{bmatrix}\left(\mathcal{H}^{D}\widetilde{x}\right)(b,t)\\ \left(\mathcal{H}^{D}\widetilde{x}\right)(a,t)\end{bmatrix},$
which allows us to e.g. straightforwardly calculate the fundamental solution [^1] Note that in the following $\Psi_{\zeta}^{s}$, $\mathbf{V}(s)$, $\mathbf{U}(s)$, $M$, etc. will always refer to the objects calculated for the diagonalised system $\textrm{pH}_{n}\left(P_{1}^{D},P_{0}^{D},\mathcal{H}_{D},\widetilde{W}_{B}^{D},\widetilde{W}_{C}^{D}\right)$. to be
(3.10) $\Psi_{\zeta}^{s}=e^{s\left(P^{D}_{1}\right)^{-1}\int_{a}^{\zeta}\left(\mathcal{H}^{D}\right)^{-1}(\xi)\,\mathrm{d}\xi}=\begin{bmatrix}e^{s\int_{a}^{\zeta}\left(\mathcal{H}^{D}_{+}\right)^{-1}(\xi)\,\mathrm{d}\xi}&0\\ 0&e^{-s\int_{a}^{\zeta}\left(\mathcal{H}^{D}_{-}\right)^{-1}(\xi)\,\mathrm{d}\xi}\end{bmatrix}.$
Note that this matrix is in particular again diagonal.
As a second step we can now also give explicit expressions for the matrices $\mathbf{V}(s)$, $\mathbf{V}(s)^{-1}$ and $\mathbf{U}(s)$ from Definition 3.8 appearing in the expression for the transfer function of the diagonalised system in Lemma 3.9,
(3.11) $\mathbf{V}(s)=\begin{bmatrix}e^{s\int_{a}^{b}\left(\mathcal{H}^{D}_{+}\right)^{-1}(\xi)\,\mathrm{d}\xi}&0\\ 0&\mathbbm{1}_{n-m}\end{bmatrix},\,\,\,\,\,\mathbf{V}(s)^{-1}=\begin{bmatrix}e^{-s\int_{a}^{b}\left(\mathcal{H}^{D}_{+}\right)^{-1}(\xi)\,\mathrm{d}\xi}&0\\ 0&\mathbbm{1}_{n-m}\end{bmatrix},$
(3.12) $\mathbf{U}(s)=\begin{bmatrix}e^{-s\int_{a}^{b}\left(\mathcal{H}^{D}_{+}\right)^{-1}(\xi)\,\mathrm{d}\xi}&0\\ 0&e^{-s\int_{a}^{b}\left(\mathcal{H}^{D}_{-}\right)^{-1}(\xi)\,\mathrm{d}\xi}\end{bmatrix}.$
Note that, while these expressions are well-defined for all $s\in\overline{\mathbb{C}_{0}}$, the expression for the transfer function found earlier is a-priori still only valid for the domain given there.
We first collect some simple observations on the matrix $\mathbf{U}(s)$.
Lemma 3.17.
Let $\mathbf{U}(s)$ be given as in Equation (3.12).
  1. (1)
    $\mathbf{U}(s)$ is a diagonal matrix with entries of the form $e^{-s\lambda_{k}}$ where $\lambda_{k}>0$.
  2. (2)
    For $s=it$, with $t\in\mathbb{R}$, the matrix $\mathbf{U}(s)$ is unitary.
  3. (3)
    For any $s\in\mathbb{C}_{0}$ the entries in $\mathbf{U}(s)$ are strictly smaller than 1.
  4. (4)
    $\|\mathbf{U}(s)\|_{\ell^{p}\rightarrow\ell^{p}}\leq 1$ for any $p\in[1,\infty)$ and $s\in\overline{\mathbb{C}_{0}}$ with equality if and only if $s\in i\mathbb{R}$.
  5. (5)
    $\|\mathbf{U}(s)\|_{\ell^{p}\rightarrow\ell^{p}}=\|\mathbf{U}({\mathrm{Re}\left(s\right)})\|_{\ell^{p}\rightarrow\ell^{p}}$ for any $p\in[1,\infty)$ and $s\in\overline{\mathbb{C}_{0}}$.
  6. (6)
    The map $t\mapsto\|\mathbf{U}(t)\|_{\ell^{p}\rightarrow\ell^{p}}$ is strictly decreasing, $t\in\mathbb{R}$, and
    $\lim_{t\rightarrow\infty}\|\mathbf{U}(t)\|_{\ell^{p}\rightarrow\ell^{p}}=0$.
  7. (7)
    $\|\mathbf{U}(s)\|_{\ell^{p}\rightarrow\ell^{p}}$ is strictly decreasing with increasing real part of $s$.
Let now $\left(M_{ij}\right)_{i,j}=M$ and consider the matrix
$M\mathbf{U}(s)=\begin{bmatrix}M_{11}e^{-s\lambda_{1}}&M_{12}e^{-s\lambda_{2}}&\cdots&M_{1n}e^{-s\lambda_{n}}\\ M_{21}e^{-s\lambda_{1}}&M_{22}e^{-s\lambda_{2}}&\cdots&M_{2n}e^{-s\lambda_{n}}\\ \vdots&\vdots&&\vdots\\ M_{n1}e^{-s\lambda_{1}}&M_{n2}e^{-s\lambda_{2}}&\cdots&M_{nn}e^{-s\lambda_{n}}\end{bmatrix},$
whose form follows from Lemma 3.17(1). Then the properties of $\mathbf{U}(s)$ from that Lemma imply convergence of the Neumann series $\sum_{k=0}^{\infty}(M\mathbf{U}(s))^{k}$ for certain $s$.
Lemma 3.18.
  1. (1)
    There exists an $\alpha_{0}\geq 0$ such that the series $\sum_{k=0}^{\infty}(M\mathbf{U}(s))^{k}$ converges for every $s\in\mathbb{C}_{\alpha_{0}}$.
  2. (2)
    If $\|M\|_{\ell^{2}\rightarrow\ell^{2}}\leq 1$ then $\sum_{k=0}^{\infty}(M\mathbf{U}(s))^{k}$ converges for every $s\in\mathbb{C}_{0}$.
  3. (3)
    If $\|M\|_{\ell^{2}\rightarrow\ell^{2}}<1$ then $\sum_{k=0}^{\infty}(M\mathbf{U}(s))^{k}$ converges for every $s\in\overline{\mathbb{C}_{0}}$.
Proof.
  1. (1)
    By Parts 5 to 7 of Lemma 3.17, there exists an $\alpha_{0}\geq 0$ such that $\|M\mathbf{U}(s)\|_{\ell^{2}\rightarrow\ell^{2}}\leq\|M\|_{\ell^{2}\rightarrow\ell^{2}}\|\mathbf{U}(s)\|_{\ell^{2}\rightarrow\ell^{2}}<1$ for $s\in\mathbb{C}_{\alpha_{0}}$. Hence the Neumann series $\sum_{k=0}^{\infty}(M\mathbf{U}(s))^{k}$ converges for these $s$.
  2. (2)
    From $\|M\|_{\ell^{2}\rightarrow\ell^{2}}\leq 1$ and $\|\mathbf{U}(s)\|_{\ell^{2}\rightarrow\ell^{2}}<1$ for $s\in\mathbb{C}_{0}$ we conclude that $\|M\mathbf{U}(s)\|_{\ell^{2}\rightarrow\ell^{2}}\leq\|\mathbf{U}(s)\|_{\ell^{2}\rightarrow\ell^{2}}<1$. Hence the Neumann series converges.
  3. (3)
    Follows directly from $\|M\mathbf{U}(s)\|_{\ell^{2}\rightarrow\ell^{2}}\leq\|M\|_{\ell^{2}\rightarrow\ell^{2}}<1$ for $\overline{\mathbb{C}_{0}}$. ∎
Using Lemma 3.18 and Lemma 3.9 we then immediately have the following results.
Corollary 3.19.
  1. (1)
    There exists $\alpha_{0}\geq 0$ such that for $s\in\mathbb{C}_{\alpha_{0}}$
    (3.13) $\mathbf{G}(s)=\mathbf{Z}(s)\sum_{k=0}^{\infty}(M\mathbf{U}(s))^{k}\,K^{-1}.$
  2. (2)
    If $\|M\|_{\ell^{2}\rightarrow\ell^{2}}\leq 1$ then (3.13) holds for $s\in\mathbb{C}_{0}$.
  3. (3)
    If $\|M\|_{\ell^{2}\rightarrow\ell^{2}}<1$ then (3.13) holds for $s\in\overline{\mathbb{C}_{0}}$.
By Lemma 3.10(2) we thus have
$\mathcal{L}^{-1}\left\{\mathbf{G}\right\}(t)=\mathcal{L}^{-1}\left\{\mathbf{Z}\right\}(t)\ast\sum_{k=0}^{\infty}\mathcal{L}^{-1}\left\{M\mathbf{U}\right\}(t)^{\ast k}K^{-1}.$
where
$\mathcal{L}^{-1}\left\{M\mathbf{U}\right\}(t)=\begin{bmatrix}M_{11}\delta(t-\lambda_{1})&M_{12}\delta(t-\lambda_{2})&\cdots&M_{1n}\delta(t-\lambda_{n})\\ M_{21}\delta(t-\lambda_{1})&M_{22}\delta(t-\lambda_{2})&\cdots&M_{2n}\delta(t-\lambda_{n})\\ \vdots&\vdots&&\vdots\\ M_{n1}\delta(t-\lambda_{1})&M_{n2}\delta(t-\lambda_{2})&\cdots&M_{nn}\delta(t-\lambda_{n})\end{bmatrix}.$
Remark 3.20.
Note in passing that, as $\mathbf{Z}\in H^{\infty}(\overline{\mathbb{C}_{\alpha}})$ and $(M\mathbf{U})^{k}\in H^{\infty}(\overline{\mathbb{C}_{\alpha}})$ for $\alpha>\alpha_{0}$, the uniform convergence implies that also $\mathbf{G}\in H^{\infty}(\overline{\mathbb{C}_{\alpha}})$.
3.5. BIBO stability of the diagonalised system
We now strive to derive sufficient conditions for BIBO stability using the condition from Part 3 of Lemma 3.10. By a slight abuse of notation aimed at conciseness, let in the following $\left\|F_{p,q}(s)\right\|_{\mathcal{M}}:=\left\|\mathcal{L}^{-1}\left\{F_{p,q}\right\}\right\|_{\mathcal{M}}$ denote the total variation of the inverse Laplace transform of the $(p,q)$-th entry of $F$.
With this notation, by Part 3 of Lemma 3.10 it is thus sufficient for BIBO stability to show that for all $1\leq p,q\leq n$ we have
$\left\|\left(\mathbf{Z}\right)_{p,q}\right\|_{\mathcal{M}}<\infty\qquad\textrm{and}\qquad\left\|\left(\left(M\mathbf{U}(s)\right)^{k}\right)_{p,q}\right\|_{\mathcal{M}}<\infty$
for any $k>0$ and that
(3.14) $\sum_{k=0}^{\infty}\left\|\left(\left(M\mathbf{U}(s)\right)^{k}\right)_{p,q}\right\|_{\mathcal{M}}<\infty.$
Lemma 3.21.
For any $\widetilde{W}_{C}\in\mathbb{K}^{n\times 2n}$ and $1\leq p,q\leq n$ we have $\left\|\left(\mathbf{Z}\right)_{p,q}\right\|_{\mathcal{M}}<\infty$.
Proof.
It follows directly from the definition of $\mathbf{Z}=\widetilde{W}_{C}\begin{bmatrix}\Psi_{b}^{s}\\ \mathbbm{1}_{n}\end{bmatrix}\mathbf{V}(s)^{-1}$ and the derived expressions for $\Psi_{b}^{s}$ and $\mathbf{V}(s)^{-1}$ that the components of $\mathbf{Z}$ are just finite linear combinations of exponential terms with negative exponent and constants. Thus their inverse Laplace transform will be a finite linear combination of delta distributions and hence a measure of bounded total variation. ∎
Considering the structure of $M\mathbf{U}(s)$ we see that any power of it will have the form
$\left(M\mathbf{U}(s)\right)^{k}=\begin{bmatrix}\sum_{i=1}^{m}\alpha^{(k)}_{11,i}e^{-s\widetilde{\lambda}_{i}}&\sum_{i=1}^{m}\alpha^{(k)}_{12,i}e^{-s\widetilde{\lambda}_{i}}&\cdots&\sum_{i=1}^{m}\alpha^{(k)}_{1n,i}e^{-s\widetilde{\lambda}_{i}}\\ \sum_{i=1}^{m}\alpha^{(k)}_{21,i}e^{-s\widetilde{\lambda}_{i}}&\sum_{i=1}^{m}\alpha^{(k)}_{22,i}e^{-s\widetilde{\lambda}_{i}}&\cdots&\sum_{i=1}^{m}\alpha^{(k)}_{2n,i}e^{-s\widetilde{\lambda}_{i}}\\ \vdots&\vdots&&\vdots\\ \sum_{i=1}^{m}\alpha^{(k)}_{n1,i}e^{-s\widetilde{\lambda}_{i}}&\sum_{i=1}^{m}\alpha^{(k)}_{n2,i}e^{-s\widetilde{\lambda}_{i}}&\cdots&\sum_{i=1}^{m}\alpha^{(k)}_{nn,i}e^{-s\widetilde{\lambda}_{i}}\end{bmatrix}$
with some coefficients $\alpha^{(k)}_{pq,i}$ and pairwise different $\widetilde{\lambda}_{i}\in\mathbb{R}$ with $1\leq i\leq m$ for some $m\leq n$ and $1\leq p,q\leq n$. Its inverse Laplace transform is then given by
$\left(M\mathbf{U}(s)\right)^{k}=\begin{bmatrix}\sum_{i}\alpha^{(k)}_{11,i}\delta\left(t-\widetilde{\lambda}_{i}\right)&\sum_{i}\alpha^{(k)}_{12,i}\delta\left(t-\widetilde{\lambda}_{i}\right)&\cdots&\sum_{i}\alpha^{(k)}_{1n,i}\delta\left(t-\widetilde{\lambda}_{i}\right)\\ \sum_{i}\alpha^{(k)}_{21,i}\delta\left(t-\widetilde{\lambda}_{i}\right)&\sum_{i}\alpha^{(k)}_{22,i}\delta\left(t-\widetilde{\lambda}_{i}\right)&\cdots&\sum_{i}\alpha^{(k)}_{2n,i}\delta\left(t-\widetilde{\lambda}_{i}\right)\\ \vdots&\vdots&&\vdots\\ \sum_{i}\alpha^{(k)}_{n1,i}\delta\left(t-\widetilde{\lambda}_{i}\right)&\sum_{i}\alpha^{(k)}_{n2,i}\delta\left(t-\widetilde{\lambda}_{i}\right)&\cdots&\sum_{i}\alpha^{(k)}_{nn,i}\delta\left(t-\widetilde{\lambda}_{i}\right)\end{bmatrix}$
and hence $\left\|\left(\left(M\mathbf{U}(s)\right)^{k}\right)_{p,q}\right\|_{\mathcal{M}}=\sum_{i=1}^{m}\left|\alpha^{(k)}_{pq,i}\right|<\infty$ with $1\leq p,q\leq n$ always has bounded total variation. For sufficient conditions for BIBO stability it thus remains to guarantee that the sums in Equation (3.14) are all finite. We do this using a number of different additional assumptions on $M$.
Theorem 3.22.
Let $\textrm{pH}_{n}\left(P^{D}_{1},0,\mathcal{H}^{D},\widetilde{W}^{D}_{B},\widetilde{W}^{D}_{C}\right)$ be a pHS in diagonalised form satisfying Assumption 3.11. Then if any of the following conditions is satisfied, Equation (3.14) holds and hence the system is BIBO stable.
  1. (1)
    $\|M\|_{\ell^{\infty}\rightarrow\ell^{\infty}}<1$
  2. (2)
    $\sum_{k=0}^{\infty}\left(M_{|\cdot|}\right)^{k}$ converges. This is in particular the case if $\|M_{|\cdot|}\|<1$ in some submultiplicative matrix norm.
  3. (3)
    There exists a $k_{0}>0$ such that $\sum_{r=1}^{n}\left\|\left(\left(M\mathbf{U}(s)\right)^{k_{0}}\right)_{p,r}\right\|_{\mathcal{M}}<1$ for $1\leq p\leq n$.
Proof.
  1. (1)
    Denote $m_{\infty}:=\|M\|_{\ell^{\infty}\rightarrow\ell^{\infty}}<1$. From the form of $M\mathbf{U}(s)$ it then follows in particular that $\sum_{r}\left\|\left(M\mathbf{U}(s)\right)_{p,r}\right\|_{\mathcal{M}}\leq m_{\infty}$ for any $p$.
    We now claim that $\sum_{r}\left\|\left(\left(M\mathbf{U}(s)\right)^{k}\right)_{p,r}\right\|_{\mathcal{M}}\leq m_{\infty}^{k}$ and prove this by induction. By what we just stated this is true for $k=1$. Let it be true for some $k$ and let $\left(M\mathbf{U}(s)\right)^{k}$ have the form as above. Then
    $\left(\left(M\mathbf{U}(s)\right)^{k+1}\right)_{p,q}$ $=\begin{bmatrix}\sum_{i}\alpha^{(k)}_{p1,i}e^{-s\widetilde{\lambda}_{i}}&\sum_{i}\alpha^{(k)}_{p2,i}e^{-s\widetilde{\lambda}_{i}}&\cdots&\sum_{i}\alpha^{(k)}_{pn,i}e^{-s\widetilde{\lambda}_{i}}\end{bmatrix}\begin{bmatrix}M_{1q}e^{-s\lambda_{q}}\\ M_{2q}e^{-s\lambda_{q}}\\ \vdots\\ M_{nq}e^{-s\lambda_{q}}\end{bmatrix}$
    $=\sum_{r}\sum_{i}\alpha^{(k)}_{pr,i}e^{-s\widetilde{\lambda}_{i}}M_{rq}e^{-s\lambda_{q}}=\sum_{r}\sum_{i}\alpha^{(k)}_{pr,i}M_{rq}e^{-s(\widetilde{\lambda}_{i}+\lambda_{q})}.$
    Thus
    $\left\|\left(\left(M\mathbf{U}(s)\right)^{k+1}\right)_{p,q}\right\|_{\mathcal{M}}\leq\sum_{i}\left|\sum_{r}\alpha^{(k)}_{pr,i}M_{rq}\right|$ $\leq\sum_{i}\sum_{r}\left|\alpha^{(k)}_{pr,i}\right|\left|M_{rq}\right|$
    $=\sum_{r}\left|M_{rq}\right|\left\|\left(\left(M\mathbf{U}(s)\right)^{k}\right)_{p,r}\right\|_{\mathcal{M}}.$
    Hence
    $\sum_{q}\left\|\left(\left(M\mathbf{U}(s)\right)^{k+1}\right)_{p,q}\right\|_{\mathcal{M}}\leq\sum_{q}\sum_{r}\left|M_{rq}\right|\left\|\left(\left(M\mathbf{U}(s)\right)^{k}\right)_{p,r}\right\|_{\mathcal{M}}$
    $\qquad=\sum_{r}\left\|\left(\left(M\mathbf{U}(s)\right)^{k}\right)_{p,r}\right\|_{\mathcal{M}}\sum_{q}\left|M_{rq}\right|\leq m_{\infty}\sum_{r}\left\|\left(\left(M\mathbf{U}(s)\right)^{k}\right)_{p,r}\right\|_{\mathcal{M}}\leq m_{\infty}^{k+1}.$
    This proves the claim and by a Neumann series argument the proposition.
  2. (2)
    For this it is sufficient to show by induction that for any $k\in\mathbb{N}$
    $\left\|\left(\left(M\mathbf{U}(s)\right)^{k}\right)_{p,q}\right\|_{\mathcal{M}}\leq\left(\left(M_{|\cdot|}\right)^{k}\right)_{p,q}.$
    The claim is clear from the expression for $M\mathbf{U}(s)$ for $k=1$. Now suppose the inequality holds for $k\in\mathbb{N}$. Then
    $\left(\left(M\mathbf{U}(s)\right)^{k+1}\right)_{p,q}$ $=\begin{bmatrix}\sum_{i}\alpha^{(k)}_{p1,i}e^{-s\widetilde{\lambda}_{i}}&\sum_{i}\alpha^{(k)}_{p2,i}e^{-s\widetilde{\lambda}_{i}}&\cdots&\sum_{i}\alpha^{(k)}_{pn,i}e^{-s\widetilde{\lambda}_{i}}\end{bmatrix}\begin{bmatrix}M_{1q}e^{-s\lambda_{q}}\\ M_{2q}e^{-s\lambda_{q}}\\ \vdots\\ M_{nq}e^{-s\lambda_{q}}\end{bmatrix}$
    $=\sum_{r}\sum_{i}\alpha^{(k)}_{pr,i}e^{-s\widetilde{\lambda}_{i}}M_{rq}e^{-s\lambda_{q}}=\sum_{r}\sum_{i}\alpha^{(k)}_{pr,i}M_{rq}e^{-s(\widetilde{\lambda}_{i}+\lambda_{q})}.$
    Thus
    $\left\|\left(\left(M\mathbf{U}(s)\right)^{k+1}\right)_{p,q}\right\|_{\mathcal{M}}\leq\sum_{r}\sum_{i}\left|\alpha^{(k)}_{pr,i}M_{rq}\right|=\sum_{r}\left|M_{rq}\right|\sum_{i}\left|\alpha^{(k)}_{pr,i}\right|$
    $\qquad=\sum_{r}\left|M_{rq}\right|\left\|\left(\left(M\mathbf{U}(s)\right)^{k}\right)_{p,q}\right\|_{\mathcal{M}}\leq\sum_{r}\left|M_{rq}\right|\left(\left(M_{|\cdot|}\right)^{k}\right)_{p,r}=\left(\left(M_{|\cdot|}\right)^{k+1}\right)_{p,r}.$
  3. (3)
    Realise that it is sufficient to show that the sum
    $\sum_{k=0}^{\infty}(M\mathbf{U}(s))^{k}=\left(\sum_{k=0}^{k_{0}-1}(M\mathbf{U}(s))^{k}\right)\left(\sum_{\ell=0}^{\infty}(M\mathbf{U}(s))^{\ell k_{0}}\right).$
    is the Laplace transform of a measure of bounded total variation. The first term is always the Laplace transform of a measure of bounded total variation as a finite sum of such terms. For the second term, let $m_{k_{0}}:=\max_{p}\sum_{r}\left\|\left(\left(M\mathbf{U}(s)\right)^{k_{0}}\right)_{p,r}\right\|_{\mathcal{M}}<1$. We prove by induction that $\sum_{r}\left\|\left(\left(M\mathbf{U}(s)\right)^{\ell k_{0}}\right)_{p,r}\right\|_{\mathcal{M}}\leq m_{k_{0}}^{\ell}$. For $\ell=1$ this is clear. For the induction consider
    $\left(\left(M\mathbf{U}(s)\right)^{(\ell+1)k_{0}}\right)_{p,q}$
    $\hskip 15.0pt=\begin{bmatrix}\sum_{i}\alpha^{(\ell k_{0})}_{p1,i}e^{-s\widetilde{\lambda}_{i}}&\sum_{i}\alpha^{(\ell k_{0})}_{p2,i}e^{-s\widetilde{\lambda}_{i}}&\cdots&\sum_{i}\alpha^{(\ell k_{0})}_{pn,i}e^{-s\widetilde{\lambda}_{i}}\end{bmatrix}\begin{bmatrix}\sum_{j}\alpha^{(k_{0})}_{1q,j}e^{-s\widetilde{\widetilde{\lambda}}_{j}}\\ \sum_{j}\alpha^{(k_{0})}_{2q,j}e^{-s\widetilde{\widetilde{\lambda}}_{j}}\\ \vdots\\ \sum_{j}\alpha^{(k_{0})}_{nq,j}e^{-s\widetilde{\widetilde{\lambda}}_{j}}\end{bmatrix}$
    $\hskip 15.0pt=\sum_{r}\left(\sum_{i}\alpha^{(\ell k_{0})}_{pr,i}e^{-s\widetilde{\lambda}_{i}}\right)\left(\sum_{j}\alpha^{(k_{0})}_{rq,j}e^{-s\widetilde{\widetilde{\lambda}}_{j}}\right)=\sum_{r}\sum_{i}\sum_{j}\alpha^{(\ell k_{0})}_{pr,i}\alpha^{(k_{0})}_{rq,j}e^{-s(\widetilde{\lambda}_{i}+\widetilde{\widetilde{\lambda}}_{j})}$
    so that
    $\left\|\left(\left(M\mathbf{U}(s)\right)^{(\ell+1)k_{0}}\right)_{p,q}\right\|_{\mathcal{M}}$
    $\qquad\leq\sum_{r}\sum_{i}\sum_{j}\left|\alpha^{(\ell k_{0})}_{pr,i}\alpha^{(k_{0})}_{rq,j}\right|=\sum_{r}\sum_{i}\left|\alpha^{(\ell k_{0})}_{pr,i}\right|\left\|\left(\left(M\mathbf{U}(s)\right)^{k_{0}}\right)_{r,q}\right\|_{\mathcal{M}}.$
    Hence we find
    $\sum_{q}\left\|\left(\left(M\mathbf{U}(s)\right)^{(\ell+1)k_{0}}\right)_{p,q}\right\|_{\mathcal{M}}$ $\leq\sum_{r}\sum_{i}\left|\alpha^{(\ell k_{0})}_{pr,i}\right|m_{k_{0}}\leq m_{k_{0}}^{\ell+1}.$
    The statement then follows again from a Neumann series argument. ∎
It is easy to see that the norm condition from part (1) of Theorem 3.22 is only a sufficient but not a necessary condition.
Example 3.23.
Consider $P_{1}=\mathcal{H}=\mathbbm{1}_{2}$ and $M=\begin{bmatrix}\frac{3}{4}&\frac{1}{2}\\ 0&\frac{1}{2}\end{bmatrix}$. Then $\|M\|_{\ell^{2}\rightarrow\ell^{2}}<1$ but $\|M\|_{\ell^{\infty}\rightarrow\ell^{\infty}}=\frac{5}{4}>1$. Explicit calculations show
$\left(M\mathbf{U}(s)\right)^{k}=\begin{bmatrix}\left(\frac{3}{4}\right)^{k}e^{-ks}&2^{1-k}\left(\left(\frac{3}{2}\right)^{k}-1\right)e^{-ks}\\ 0&\left(\frac{1}{2}\right)^{k}e^{-ks}\end{bmatrix}$
and hence $\mathcal{L}^{-1}\left\{\left(M\mathbf{U}\right)^{k}\right\}(t)=\begin{bmatrix}\left(\frac{3}{4}\right)^{k}\delta(t-k)&2^{1-k}\left(\left(\frac{3}{2}\right)^{k}-1\right)\delta(t-k)\\ 0&\left(\frac{1}{2}\right)^{k}\delta(t-k)\end{bmatrix}.$
As $\sum_{k=0}^{\infty}\left(\frac{3}{4}\right)^{k}<\infty$, $\sum_{k=0}^{\infty}\left(\frac{1}{2}\right)^{k}<\infty$ and $\sum_{k=0}^{\infty}2^{1-k}\left(\left(\frac{3}{2}\right)^{k}-1\right)<\infty$, we can conclude that $\mathcal{L}^{-1}\left\{\mathbf{G}\right\}$ is a measure of bounded total variation and hence the system is BIBO stable.
Furthermore, the following example provides a case where neither the sufficient condition from part (1) nor the one from part (2) of Theorem 3.22 are satisfied, but where part (3) can be applied.
Example 3.24.
Consider $\textrm{pH}_{2}\left(\mathbbm{1}_{2},0,2\mathbbm{1}_{2},\left[\begin{smallmatrix}1&0&\frac{1}{2}&\frac{1}{2}\\ 0&1&\frac{1}{2}&-\frac{1}{2}\end{smallmatrix}\right],\left[\begin{smallmatrix}0&0&1&0\\ 0&0&0&1\end{smallmatrix}\right]\right)$. Then we find
$\mathbf{G}(s)=\begin{bmatrix}\frac{1-2e^{\frac{s}{2}}}{1-2e^{s}}&\frac{1}{1-2e^{s}}\\ \frac{1}{1-2e^{s}}&-\frac{1+2e^{\frac{s}{2}}}{1-2e^{s}}\end{bmatrix}$ for which we cannot easily find the inverse Laplace transform and use it to check for BIBO stability.
Now employing the decomposition from Proposition 3.13, we find $M=\frac{1}{2}\left[\begin{smallmatrix}-1&-1\\ -1&1\end{smallmatrix}\right]$. As $\|M\|_{\ell^{\infty}\rightarrow\ell^{\infty}}=1$ we cannot apply part (1) and as $\left(M_{|\cdot|}\right)^{k}=\frac{1}{2}\left[\begin{smallmatrix}1&1\\ 1&1\end{smallmatrix}\right]$ the sum $\sum_{k=0}^{\infty}\left(M_{|\cdot|}\right)^{k}$ diverges thus ruling out the use of part (2) of 3.22.
However, we find that $(M\mathbf{U}(s))^{2}=\frac{e^{s}}{2}\left[\begin{smallmatrix}1&0\\ 0&1\end{smallmatrix}\right]$ and thus
$\left\|\mathcal{L}^{-1}\left\{\left(M\mathbf{U}\right)^{2}\right\}(t)\right\|_{\mathcal{M}}=\left\|\begin{bmatrix}\frac{\delta(t)}{2}&0\\ 0&\frac{\delta(t)}{2}\end{bmatrix}\right\|_{\mathcal{M}}=\begin{bmatrix}\frac{1}{2}&0\\ 0&\frac{1}{2}\end{bmatrix}$
implying by Theorem 3.22(3) that the pHS is BIBO stable.
4. Another approach to BIBO
Another approach to BIBO stability of pHS that we want to present here is based upon the $L^{1}$-well-posedness of these systems when considered on the state space $L^{1}$.
Lemma 4.1.
A pHS satisfying Assumptions 2.1 and 3.11 is $L^{1}$-well-posed if considered as a system on the state space $L^{1}$, that is for some (and hence all) $t>0$ there exists a constant $c>0$ such that for any classical solution $(u,x,y)$ we have that
$\|y\|_{L^{1}\left([0,t],Y\right)}+\|x(t)\|_{L^{1}\left([a,b],\mathbb{K}\right)}\leq c\left(\|u\|_{L^{1}\left([0,t],U\right)}+\|x(0)\|_{L^{1}\left([a,b],\mathbb{K}\right)}\right).$
Proof.
By Assumption 2.1 and [11, Thm. 1.5] the operator $A$ considered as defined analogously on a domain in $L^{1}\left([a,b],\mathbb{K}^{n}\right)$ generates a $C_{0}$-semigroup on $L^{1}\left([a,b],\mathbb{K}^{n}\right)$ and thus by [35, Thm. 7.1] the pHS considered on the state space $L^{1}\left([a,b],\mathbb{K}^{n}\right)$ is $L^{1}$-well-posed. ∎
Remark 4.2.
  1. (1)
    Note that while [11, Thm. 1.5] is formulated only for the case of a spatially constant $P_{0}$ the proof thereof does not make use of this property and it thus also holds in the more general case considered here.
  2. (2)
    It is essential here to consider the state space $L^{1}\left([a,b],\mathbb{K}^{n}\right)$, as on any reflexive Banach space (in particular any $L^{p}$ with $1<p<\infty$) $L^{1}$-well-posedness already implies boundedness of the input operator [23, Thm. 4.2.7].
Theorem 4.3.
If a port-Hamiltonian system satisfying Assumptions 2.1 and 3.11 is exponentially stable on the state space $\widetilde{X}=L^{1}\left([a,b],\mathbb{K}^{n}\right)$ then it is BIBO stable.
Proof.
By Lemma 4.1 we have that the pHS is $L^{1}$-well-posed on $\widetilde{X}$. But then by [23, Lem. 4.5.1] respectively an argument similar to [3, Lem. 3.22] in the general case of non-compatible systems we can show that the system is $C^{\infty}$-LILO stable in the sense of [20, Def. 4.1] and thus by [20, Thm. 4.2 & Cor. 4.4] also BIBO stable. ∎
While this theorem provides us with a sufficient condition for BIBO stability it leaves us with the problem of having to determine whether the semigroup generated by the operator $A$ on $L^{1}$ is exponentially stable. While it is easy to e.g. show that the spectrum of the operator $A$ considered on $L^{1}$ agrees with the one it has on $L^{2}$, it seems unclear so far which connection (if any) there is between exponential stability on $L^{2}$ and exponential stability on $L^{1}$.
Furthermore, many approaches to checking for exponential stability, such as the Gearhart-Prüss theorem — which forms the basis of the study of exponential stability of pHS in [24] — cannot be applied in the $L^{1}$ case as they only hold on Hilbert spaces.
Results of a general nature implying exponential stability of a system on $L^{1}$ seem to exist only in restricted cases such as e.g. for positive semigoups [16, Thm. 3.5.3].
5. Conclusion
5.1. Some conjectures
During our studies of BIBO stability for distributed port-Hamiltonian systems underlying this work, a few observations have been made that suggest certain conjectures.
First, we have so far not encountered a system that satisfies the condition (2.2) as a strict inequality and is not BIBO stable. Considering how (2.2) relates to the norm of $M$ and the role this matrix plays in the construction of the counterexample in Proposition 2.5 as well as in the arguments in [24, Thm. 5.6], motivates us to propose the following conjecture on a sufficient condition for BIBO stability.
Conjecture 5.1.
If $\widetilde{W}_{B}R_{0}^{-1}\Sigma\left(\widetilde{W}_{B}R_{0}^{-1}\right)^{*}>0$ the pHS is BIBO stable.
Note that by Remark 3.4 this condition is — in the case that Assumption 3.5 holds — equivalent to $\|M\|_{\ell^{2}\rightarrow\ell^{2}}<1$.
That the condition from this conjecture, could only be a sufficient condition for BIBO stability follows from the following example, which is based upon a class of systems discussed in other contexts in [7, Sec. 5] and [24, Ex. 4.4].
Example 5.2.
Consider the system $\textrm{pH}_{2}\left(\mathbbm{1}_{2},0,\mathbbm{1}_{2},\widetilde{W}_{B},\widetilde{W}_{C}\right)$ shown to be BIBO stable in the proof of Proposition 2.6, but for which $\widetilde{W}_{B}R_{0}^{-1}\Sigma\left(\widetilde{W}_{B}R_{0}^{-1}\right)^{*}\ngtr 0$.
One potential way to approach Conjecture 5.1 — at least in the case of the diagonalised system and $P_{0}^{D}=0$ — may be via Theorem 3.22(3). Indeed, we have not yet found an example for a system with $\|M\|_{\ell^{2}\rightarrow\ell^{2}}<1$ for which no $k_{0}$ as required in this proposition exists. This motivates us to propose the following conjecture which would imply Conjecture 5.1 for this subclass of pHS.
Conjecture 5.3.
Let $\|M\|_{\ell^{2}\rightarrow\ell^{2}}<1$. Then there exists a $k_{0}$ such that
$\sum_{r}\left\|\left(\left(M\mathbf{U}(s)\right)^{k_{0}}\right)_{p,r}\right\|_{\mathcal{M}}<1,\qquad 1\leq p\leq n.$
Note that the strict inequality is necessary here, as for the case $\|M\|_{\ell^{2}\rightarrow\ell^{2}}=1$ there are counterexamples.
Example 5.4.
[…]
$\left(M\mathbf{U}(s)\right)^{k}=\frac{(-1)^{k}}{2^{k}}\sum_{i=0}^{k-1}\binom{k-1}{i}e^{-s(2k-i)}(-1)^{i}\begin{bmatrix}1&e^{s}\\ -1&-e^{s}\end{bmatrix}.$
Therefore we find $\left\|\left(M\mathbf{U}(s)\right)^{k}\right\|_{\mathcal{M}}=\frac{1}{2^{k}}\sum_{i=0}^{k-1}\binom{k-1}{i}\begin{bmatrix}1&1\\ 1&1\end{bmatrix}=\frac{1}{2}\begin{bmatrix}1&1\\ 1&1\end{bmatrix},$ so that
$\sum_{r}\left\|\left(\left(M\mathbf{U}(s)\right)^{k}\right)_{p,r}\right\|_{\mathcal{M}}=1$ for all $k$.
Similar, we have so far not even yet encountered an exponentially stable pHS that is not also BIBO stable. This may motivate one to suggest the following — due to [12, Lem. 9.1.4] in some sense more general — conjecture.
Conjecture 5.5.
Any exponentially stable pHS is BIBO stable.
5.2. Outlook
While the final result of the approach discussed in Section 3 only provides conditions for a restricted case $P^{D}_{0}=0$ and thus in particular for the case of $P_{0}=0$ and a constant Hamiltonian, one could hope that these can be straightforwardly extended to the more general case using simple perturbation arguments as is e.g. done in [12] to show $L^{2}$-well-posedness for pHS or in [21] to show BIBO stability of certain parabolic systems. Unfortunately there are convincing reasons that this does not work analogously in this case, such as the observation that the simple shift-semigroup with identity input is not BIBO stable [20, Ex. 3.1].
Besides the extension of the results to the case $P^{D}_{0}\neq 0$ there are two further evident routes for possible generalisation. The first would be to consider higher order pHS such as they were for example investigated for asymptotic stability in [32] employing an approach also based upon an adapted concept of an associated fundamental solution. The second lies in the extension to pHS of higher spatial dimensions as were studied e.g. in [29, 22]. Here, however, one has to note that the input and output spaces may be infinite-dimensional — as with boundary control along a one-dimensional boundary for instance. In this case the characterisation of BIBO stability by the transfer function employed in the first approach is no longer valid. Hence in this case only the second approach using $L^{1}$-well-posedness may lead to sufficient conditions.
Acknowledgements
The authors would like to thank Prof. M. Waurick (TU Freiberg) for helpful discussions on this subject.